if F(x)=-3x+2, what are the values of m and b so that F(mx=b)=2x-3?
Answer by adamchapman(301)

F(x)=-3x+2
Thuis means that to find F(x), you must multiply what is in the brackets on the left hand side (x) by -3, and then add 2.
F(mx+b)=-3(mx+b)+2
We know that this is equal to 2x-3, so:
-3(mx+b)=2x-3
-3mx-3b=2x-3
Equating coefficients of x:
-3m=2
m=-2/3
Equating constant terms:
-3b=-3
b=1
so the solution is m=-2/3 and b=1
F(-2x/3+1)=2x-3
I hope this helps.
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk
Adam